OpenStudy (anonymous):

John made a rectangle pen for his dog using 28 feet of fencing. If the width of the pen is 2 feet more than one-half the length, what is the length and width of the pen? The length of the pen is i got 6 feet??

OpenStudy (saifoo.khan):

Total = 28 feet W = 0.5x + 2 L = x 2(x) + 2(0.5x + 2) = 28 2x + x + 4 = 28 3x = 24 x = 8 L = 8 feet W = 6 feet Double check - (8 + 8 + 6 + 6 = 28)

OpenStudy (anonymous):

8 feet??

OpenStudy (saifoo.khan):

L = Length.

OpenStudy (anonymous):

yeah so 8 feet right?

OpenStudy (saifoo.khan):

L = Lol

OpenStudy (saifoo.khan):

Y= yess.